OpenStudy (anonymous):

if z-1/z+1 is purely imaginary number (z±-1),find the value of |z|

OpenStudy (anonymous):

(z±√-1) like this ?

OpenStudy (anonymous):

igtg, and idk this ........

OpenStudy (anonymous):

pls describe this

OpenStudy (anonymous):

Let \(z=a+bi\), then \[\begin{align*}\frac{z-1}{z+1}&=\frac{(a-1)+bi}{(a+1)+bi}\cdot\frac{(a+1)-bi}{(a+1)-bi}\\&=\frac{a^2-1+b^2}{(a+1)^2+b^2}\\&=\frac{a^2+b^2-1}{a^2+2a+1+b^2} \end{align*}\] Notice that there are no factors of \(i\) here, which means we've reduced it to a real number. But \(\dfrac{z-1}{z+1}\) is purely imaginary, which means this number must be 0. \[\frac{a^2+b^2-1}{a^2+2a+1+b^2}=0~~\Rightarrow~~a^2+b^2=1~~\Rightarrow~~|z|=\sqrt{a^2+b^2}=\sqrt1=1\]

OpenStudy (loser66):

@SithsAndGiggles I think there is mistake from numerator :) it should be a^2 +b^2+2bi-1

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

@Loser66 Yes, you're right, I made a mistake with the computation, but it still works out nicely. The point is, we want the real part of the numerator to be zero; I just happened to leave to leave out the imaginary part. @ssw, the answer is still correct, there was just a slight computational error that does not change the end result.